# Bias for Action: Video Implicit Neural Representations with Bias Modulation

### v1

@@ 2 Prior Work
We provide a brief overview of implicit neural representations (INRs) as they applied to video representations.
@@
A key component affecting INRs capacity is its non-linear activation function. The popular activation choice, ReLU, is however ill-suited for INRs due to their spectral bias towards lower frequencies [30]. Numerous works have addressed this issue, including positional encoding [26], sinusoidal activation function [38], Gaussian activation function [31], wavelet activation function [37] and many more.
@@ 3 ActINR
We now present our technique named ActINR (pronounced Actioner). We first motivate our approach by looking at the internal working of an INR, and then describe our approach.
INRs and Basis Representation. An Implicit Neural Representation (INR) provides a continuous approximation of a target signal $I:\mathbf{x}\in\mathbb{R}^{D}\rightarrow g(\mathbf{x})\in\mathbb{R}^{C}$ by utilizing a neural network $f_{\theta}:\mathbf{x}\in\mathbb{R}^{D}\rightarrow f(\mathbf{x})\in\mathbb{R}^{C}$ composed of linear layers and nonlinear activation functions. The learnable parameter set, including weight matricies and bias vectors, is collectively represented by $\theta=\{\mathbf{W}^{1},\mathbf{b}^{1},\ldots,\mathbf{W}^{L},\mathbf{b}^{L}\}$. INR is trained to represent the target signal by minimizing the mean-squared error objective between $\mathbf{g}(\mathbf{x}_{s})$ and $\mathbf{f}(\mathbf{x}_{s})$, where $\mathbf{x}_{s}$ denotes sampled input-output pairs. Specifically,
@@
Bias-motion Interplay. Our study is grounded in the hypothesis that motions in video can be modeled via the movement of compact local basis functions in an INR, governed by the INR’s bias parameters. Furthermore, assuming scene elements are largely the same in appearance over time, the weights of the INR may be shared across time. Fig. 3 illustrates these idea for a toy video in which a Gaussian blob on the left moves toward a stationary one on the right. While the morphology of the basis functions stay constant over time, basis function #1 translates rightward in response to the rightward motion of the left-hand Gaussian by changing bias values, while basis function #2 remains fixed. Basis functions #1 and #2 also minimally interfere with each other due to their bounded spatial support, facilitating local motion modeling.
@@ 4 Experiments
Implementation Details. We employ a three-layer MLP with WIRE activation functions, each hidden layer having a dimension size of 36 per each spatial block. The bias-INR hypernetwork utilizes Random Fourier Features (RFF) with a size of 40, where frequencies are sampled from a zero-mean Gaussian distribution with a variance of 5. For interpolation tasks, the scale and frequency parameters of WIRE are set to 10 and 100, respectively, while for inpainting tasks, they are set to 5 and 50. The video volume is divided into 96×96 patches, each spanning 10 frames, with a distinct MLP unit assigned to each patch. Each MLP unit is trained for 2000 iterations using the Adam optimizer with a learning rate of 5×$10^{-3}$, coupled with a step decay rate scheduler with a decay ratio of 0.1. Model optimization is carried out by minimizing the mean squared error (MSE) loss between the ground truth and predicted frames. For interpolation and denoising tasks, the model comprises 3.04 million parameters for the DAVIS dataset and 3.17 million parameters for the UVG dataset. All experiments are conducted on NVIDIA RTX 4090 GPU with 24 GB of memory.
@@ 4.1 Video processing with ActINR
Extreme interpolation. To further substantiate our findings, we devised more challenging experimental setups, skipping ten frames instead of every other frame in the training videos, including the “lava” and “child” sequence. The qualitative results for this extreme frame-skipping scenario, presented in Fig. 5, demonstrate that as the interval between training samples increases, the performance discrepancy between our method and the baselines becomes more pronounced in favor of our approach. We argue that the interaction between the basis functions and the final reconstruction serves as a form of regularization, enabling more precise motion tracking compared to optical flow-based methods, which are prone to errors. While H-NeRV Boost exhibits ghosting artifacts and FF-NeRV fails to resolve motion, resulting in pixel misalignments, our method preserves finer details as illustrated in the final column of Fig. 5, despite the extreme frame skipping.
@@ 4.2 Ablation Study
We now explore some of the design choices for our approach, wich specific emphasis on activation function, and interpolation module. Additional ablation studies are provided in the supplementary document.
@@ 5 Discussions and Conclusion
Limitation and Future Works. Our approach assumes that motion remains confined within a designated block size. Consequently, selecting a block size that allows the object within to exit its boundaries results in suboptimal performance. To address this limitation, when motion leads to objects overflowing beyond their blocks, it becomes necessary to increase the parameter size. This adjustment requires scaling the parameter size proportionally to the magnitude of the motion. Nonetheless, aligning with our objective to model local and minor motions, we utilize smaller block sizes accompanied by a relatively small number of parameters. By using more blocks with fewer parameters, our approach stabilizes the parameter count at a lower threshold, mitigating the need for more parameters to handle larger motions. Future work will explore sharing parameters across blocks to further address parameter scaling.
@@
Failure case. In challenging scenarios, we observe that transition regions across patches are not well reconstructed when objects cross patch boundaries. For example, as shown in Fig. 13, the horse’s foreleg remains on the boundary during its motion. When a patch boundary splits an object into separate parts, the model may struggle to reconstruct these regions in a visually plausible manner.

### v2

@@ 2 Prior Work
We provide a brief overview of implicit neural representations (INRs) as they are applied to video representations.
@@
A key component affecting INRs capacity is its non-linear activation function. The popular activation choice, ReLU, is however ill-suited for INRs due to their spectral bias towards lower frequencies [30]. Numerous works have addressed this issue, including positional encoding [26], sinusoidal activation function [38], Gaussian activation function [31], wavelet activation function [37], and many more.
@@ 3 ActINR
We now present our technique named ActINR. We first motivate our approach by looking at the internal working of an INR, and then describe our approach.
INRs and Basis Representation. An Implicit Neural Representation (INR) provides a continuous approximation of a target signal $I:\mathbf{x}\in\mathbb{R}^{D}\rightarrow g(\mathbf{x})\in\mathbb{R}^{C}$ by utilizing a neural network $f_{\theta}:\mathbf{x}\in\mathbb{R}^{D}\rightarrow f(\mathbf{x})\in\mathbb{R}^{C}$ composed of linear layers and nonlinear activation functions. The learnable parameter set, including weight matrices and bias vectors, is collectively represented by $\theta=\{\mathbf{W}^{1},\mathbf{b}^{1},\ldots,\mathbf{W}^{L},\mathbf{b}^{L}\}$. INR is trained to represent the target signal by minimizing the mean-squared error objective between $\mathbf{g}(\mathbf{x}_{s})$ and $\mathbf{f}(\mathbf{x}_{s})$, where $\mathbf{x}_{s}$ denotes sampled input-output pairs. Specifically,
@@
Bias-motion Interplay. Our study is grounded in the hypothesis that motions in video can be modeled via the movement of compact local basis functions in an INR, governed by the INR’s bias parameters. Furthermore, assuming scene elements are largely the same in appearance over time, the weights of the INR may be shared across time. Fig. 3 illustrates these ideas for a toy video in which a Gaussian blob on the left moves toward a stationary one on the right. While the morphology of the basis functions stays constant over time, basis function #1 translates rightward in response to the rightward motion of the left-hand Gaussian by changing bias values, while basis function #2 remains fixed. Basis functions #1 and #2 also minimally interfere with each other due to their bounded spatial support, facilitating local motion modeling.
@@ 4 Experiments
Implementation Details. We employ a three-layer MLP with WIRE activation functions, each hidden layer having a dimension size of 36 per each spatial block. The bias-INR hypernetwork utilizes Random Fourier Features (RFF) with a size of 40, where frequencies are sampled from a zero-mean Gaussian distribution with a variance of 5. For interpolation tasks, the scale and frequency parameters of WIRE are set to 10 and 100, respectively, while for inpainting tasks, they are set to 5 and 50. The video volume is divided into 96×96 patches, each spanning 10 frames, with a distinct MLP unit assigned to each patch. Each MLP unit is trained for 2000 iterations using the Adam optimizer with a learning rate of 5×$10^{-3}$, coupled with a step decay rate scheduler with a decay ratio of 0.1. Model optimization is carried out by minimizing the mean squared error (MSE) loss between the ground truth and predicted frames. For interpolation and denoising tasks, the model comprises 3.04 million parameters for the DAVIS dataset and 3.17 million parameters for the UVG dataset. All experiments are conducted on NVIDIA RTX 4090 GPU with 24 GB of memory. For more details, please refer to the supplementary material.
@@ 4.1 Video processing with ActINR
Extreme interpolation. To further substantiate our findings, we devised more challenging experimental setups, skipping ten frames instead of every other frame in the training videos, including the “lava” and “child” sequences. The qualitative results for this extreme frame-skipping scenario, presented in Fig. 5, demonstrate that as the interval between training samples increases, the performance discrepancy between our method and the baselines becomes more pronounced in favor of our approach. We argue that the interaction between the basis functions and the final reconstruction serves as a form of regularization, enabling more precise motion tracking compared to optical flow-based methods, which are prone to errors. While H-NeRV Boost exhibits ghosting artifacts and FF-NeRV fails to resolve motion, resulting in pixel misalignments, our method preserves finer details as illustrated in the final column of Fig. 5, despite the extreme frame skipping.
@@ 4.2 Ablation Study
We now explore some of the design choices for our approach, with specific emphasis on activation function, and interpolation module. Additional ablation studies are provided in the supplementary document.
@@ 5 Discussions and Conclusion
Limitation and Future Works. Our approach assumes that motion remains confined within a designated block size. Consequently, choosing a block that allows objects to exit its boundaries leads to suboptimal performance, and accommodating overflow requires increasing the parameter size proportionally to the magnitude of the motion. Nonetheless, by targeting local and minor motions, we rely on smaller blocks with fewer parameters, which stabilizes the overall parameter count and mitigates the need for larger models. Although our primary focus is on investigating the relationship between biases and motion, we also conducted compression experiments to illustrate performance trade-offs. In these tests, our method performs slightly worse than Hybrid NeRV while outperforming JPEG2000 (see the supplementary for details). Future work will explore sharing parameters across blocks to further address parameter scaling.
@@
Failure case. In challenging scenarios such as Jockey, we observe that transition regions across patches are not well reconstructed when objects cross patch boundaries. For example, as shown in Fig. 13, the horse’s foreleg remains on the boundary between two patches during its motion. When a patch boundary splits an object into separate parts, the model may struggle to reconstruct these regions in a visually plausible manner.
@@
Resilience to higher resolution. We evaluate the resolution invariance of our interpolation method on four 4K-resolution videos from the UVG dataset. As presented in Tab. 6, our method maintains image quality even as the video resolution increases.
Video Yacht Ready Jockey Bosph
ActINR-4K 30.1 25.7 24.0 37.2
ActINR-HD 30.3 25.9 24.1 37.3
Table 6: Interpolation performance on 4K and HD resolution
Method Encoding Time Decoding FPS
ActINR 5h19m 59.28
FF-NeRV 6h0m 84.00
HNeRV-Boost 1h53m 13.15
Table 7: Per-video encoding speed, and decoding performance
Further motivation for bias-motion interplay. To emphasize the relationship between biases and motion, we took video of a vibrating tuning fork, and fit our ActINR to it. We then median filtered the time series of biases across time to smoothen the motion artifacts, as shown in Fig. 15. As evident, the high frequency parts around the fork are replaced by a near-static prongs, thereby underscoring the relationship between biases and motions.
Encoding and Decoding speeds. Table. 7 tabulates encoding and decoding times for various approaches. Our encoding and decoding times are comparable to previous approaches, with a performance that is similar to FF-NeRV while decoding 4.5× faster than HNeRV-Boost and operating in real time, unlike HNeRV-Boost. We achieve speed up by evaluating all windows in a frame at a time (pytorch bmm). We also attribute the speed-up to initializing each group with the preceding group’s weights, reducing the training epochs needed for convergence.
Further implementation details. We reproduced FF-NeRV using a 3-million-parameter model for a fair comparison, since the original FF-NeRV reported PSNR only for its 12-million-parameter version. Following the official repository, we kept all hyperparameters unchanged except for model size. Our approach consistently employs a 3-million-parameter model for both interpolation and denoising tasks. However, for inpainting, we used triple the capacity to handle missing regions more effectively.
Figure 15: Video of a vibrating tuning fork. Upon applying a median filter to the biases, the pronounced shaking is significantly mitigated, as evident in the image on the right.
Figure 16: Rate-Distortion curve for Bosphorus
Compression. For a fairer comparison, we evaluated the compression performance of Act-INR with simple pruning and 8-bit post-quantization. Fig. 16 shows a rate distortion curve illustrating that Act-INR performs considerably better than JPEG2000, although, slightly worse than Hybrid NeRV. These analyses demonstrate that ActINR primarily excels in solving inverse problems (interpolation, super resolution, denoising, and inpainting) while state-of-the-art techniques like Hybrid NeRV [4] are better suited for compression tasks.